Acme Tools manufactures and sells cordless in a market where price (p) and demand (D) are related as follows:

p = $40 + (3,000)/D – (4,800)/D2

 

The fixed cost (CF) is $1000 per month and the variable cost per drill (cv) is $46.

 

a)     How many drills should be produced each month to maximize profits?

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

b)     How much is the maximum profit each month?

 

 


1
Expert's answer
2021-05-20T11:36:34-0400

Given;

"P = \$40 + \frac{3,000}{D} – \frac{4,800}{D^2}.................................(i)"


"FC=\$1000" and "VC=\$46" per drill, "TC=\$1000+46D"


(a) Number of drills to be produced each month to maximize profits

Profit is maximized at the point where MR=MC

"TR=P\times D"

"=[40+\frac{3000}{D}-\frac{4800}{D^2}]\times D"


"=40D+3000-\frac{4800}{D}"


"MR=40-(-\frac{4800}{D^2})"


"=40+\frac{4800}{D^2}"


TC=$1000+46D


"MC=46"


Taking MR=MC,

"40+\frac{4800}{D^2}=46"


"\frac{4800}{D^2}=46-40"


"\frac{4800}{D^2}=6"


"4800=6D^2"


"800=D^2"


"D=28.2843\approx28"


(b)

"Profit=TR-TC"

Substituting D in equation (i) we calculate the price as follows:

"P = \$40 + \frac{3,000}{28} – \frac{4,800}{28^2}"


"P=141.02"


"Profit=(P\times D)-(TC)"

"=(141.02\times 28)-[1000+46(28)]\\3,948.56-2,288=1,660.56"


Therefore the maximum profit per month is $1,660.56